OpenStudy (anonymous):

The tax on a new car costing $14,400 is $784. At the same tax rate, what is the tax on a vehicle costing $20,000? Round to the nearest dollar. $1,089 $1,568 $1,278 $989

OpenStudy (anonymous):

The answer is (A) 784/14,400 = x/20,000 show the works: 784/14,400 = x/20,000 (20000 * 784) = 14400x x = 20000/14400 = x/784 x = (20000 * 784)/14400 x = $1089 The answer is (A) $1,089 hope this help☻